OpenStudy (anonymous):

What is csc(pi/3)? Round to the nearest thousandth if necessary.

OpenStudy (anonymous):

1.1547

OpenStudy (anonymous):

How did you get that?

OpenStudy (ranga):

\[\Large \csc(\pi /3) = \frac{1}{\sin(\pi /3)} = \frac{1}{\sqrt{3}/2} = \frac{2}{\sqrt{3}} = ?\]